# Get Your Embedding Space in Order: Domain-Adaptive Regression for Forest Monitoring

### v1

@@ Abstract
Image-level regression is an important task in Earth observation, where visual domain and label shifts are a core challenge hampering generalization. However, cross-domain regression with remote sensing data remains understudied due to the absence of suited datasets. We introduce a new dataset with aerial and satellite imagery in five countries with three forest-related regression tasks. To match real-world applicative interests, we compare methods through a restrictive setup where no prior on the target domain is available during training, and models are adapted with limited information during testing. Building on the assumption that ordered relationships generalize better, we propose manifold diffusion for regression as a strong baseline for transduction in low-data regimes. Our comparison highlights the comparative advantages of inductive and transductive methods in cross-domain regression.
@@ 1 Introduction
Data availability and downstream applications define the scope of models (regional [29] - national [36] - continental [5]), but they don’t necessarily match the visual and semantic domains defined by biomes and environmental conditions. Intuitively, a model trained on a dataset should perform relatively well on datasets that are spatially close and visually similar, although it remains unclear to what extent.
To better assess generalization performance and go towards universal models that can be trained in countries with high data availability and applied to any country without prior knowledge or retraining, we introduce the DRIFT [^1] Dataset will be made publicly available upon publishing dataset for universal, low-data domain-adaptive regression.
Building on recent advances in order learning, we hypothesize that order relationships generalize well across domains. We enforce a well-ordered embedding space in the source domain, then adapt to target domains in low-data setups. Notably, we show that this strong regularization during training does not hamper source-domain performance and creates a clean embedding space to propagate similarities at test-time, which opens the way for advanced low-data inference in the form of manifold diffusion. Our contributions are as follows:
@@
  • We compare inductive and transductive methods across datasets and tasks, in a low-shot universal and source-free domain adaptation setup. We show that transductive methods take advantage when the domain gap becomes predominant.
@@ 2.3 Image-level regression in Earth observation
The increasing availability of vast and diverse collections of remote sensing imagery and datasets presents both opportunities and challenges for computer vision tasks. To date, most works utilizing remote sensing imagery focus on image classification and semantic segmentation [2, 40, 4]. For example, to count trees, some works rely on the detection of individual trees [26, 5], but manual annotation relies on expert knowledge and is extremely time-consuming. Given the variations in ground resolution, image quality, and geographic areas, small objects and occlusion phenomena can make exhaustive annotation virtually impossible. Other labels on woody biomass or canopy height cannot be manually labeled. Therefore, labels often have to be sourced from field measures [30, 9], or derived from low-resolution dedicated sensors such as GEDI [22, 13], and remain noisy due to the varying viewing angles, technical inaccuracies in positioning systems [15] or temporal mismatch [45, 6]. As such, accurate labels are often found at a lower resolution than the ground resolution of the image. This makes image-level regression a core task in computer vision for Earth observation, as it offers a way to directly predict the information at the available resolution [12, 7]. To our knowledge, there is no existing dataset for cross-domain image-level regression using sub-meter resolution remote sensing imagery.
@@ 3 DRIFT: Domain-adaptive Regression for Forest Monitoring across Countries
    Tree count: number of overstory trees in the images.
@@
Dataset collection. We built the dataset from national aerial orthophotography with various ground sampling distances (GSD) for Denmark, Spain, France, and Slovakia. All aerial images contain three spectral channels: Red (R), Green (G), and Blue (B). For Slovenia, we used very high-resolution satellite imagery from commercial nanosatellites (SkySat). We acquired the orthorectified surface reflectance products at a spatial resolution of 50cm.
Table 1: Dataset characteristics. GSD and Height in meters, Cover in percents.
@@
Figure 2: Challenging examples in the DRIFT dataset: despite similar visual content, images can have different values in the label space. More visual examples can be found in Supplementary Fig. 1
@@
Figure 3: Label shift across countries. Wasserstein distances between label distributions in country subsets indicate the level of label shift. Note that the shifts depend on the target variable.
Regression tasks. Canopy height, tree count, and tree cover heavily depend on tree species, local climate, and landscape characteristics. Height prediction from optical images is inherently challenging due to limited visual clues about depth from an aerial viewpoint, especially if the ground is not visible, but it has been proven possible using satellite imagery [22, 33]. Counts depend on the model’s interpretation of individual crown area, which is subject to extreme variations even within a species [18]. Cover fraction is intuitively the easiest, as the model only needs to separate trees from background pixels. The difficulty depends on the level of similarity between short vegetation ($<5$m) and tree cover.
@@ 4 Method
We propose to explore the potential of ordered embeddings and simple transductive learning on the cross-domain challenges of the DRIFT dataset (see Fig. 4). We train the embedding extractor on the source domain, using geometric order learning to enforce structure. We extract embeddings on the target domain, and perform adaptation for regression using a few labeled images. Optionally, we refine predictions to take into account local and global manifold structure with diffusion.
@@ 4.1 Preliminaries
Let $\mathcal{D}_{s}{=}\{(\textbf{x}_{i}^{s},y_{i}^{s})\}$ and $\mathcal{D}_{t}{=}\{(\textbf{x}_{i}^{t},y_{i}^{t})\}$ denote the source and target domains respectively. $y^{s}$ and $y^{t}$ take continuous values in $\mathbb{R}$, and we note that their distributions are different. Notably, it is common in vegetation mapping that their support $[y^{s}_{min},y^{s}_{max}]$ and $[y^{t}_{min},y^{t}_{max}]$ differs. For example, $y^{s}_{max}\neq y^{t}_{max}$ in height mapping due to different species.
@@
A simple baseline method would be to train a regressor $f(\textbf{x})=y$ on $\mathcal{D}_{s}$ and apply it to $\mathcal{D}_{t}$. Due to domain shifts, this performs generally poorly. Instead, we consider a feature extractor $h(\textbf{x})=\mathbf{v}$, with $\mathbf{v}\in\mathbb{R}^{d}$ and $\lVert\mathbf{v}\lVert_{2}=1$, and frame the problem as cross-domain transduction: $y^{t}$ is inferred from the relationships between the embeddings in $\mathcal{D}_{t}$.
@@ 4.2 Manifold Diffusion for Regression
We start by introducing the diffusion framework originally proposed in [51]. Assuming a set $X=\{\textbf{x}_{1},\textbf{x}_{2},...,\textbf{x}_{n}\}$ and $\mathbf{v}_{i}=h(\textbf{x}_{i})$, the affinity matrix $W\in\mathbb{R}^{n\times n}$ is constructed by calculating the pairwise cosine similarity between samples:
$a_{ij}:=\begin{cases}[\mathbf{v}_{i}^{\top}\mathbf{v}_{j}]_{+}^{\gamma},&\textrm{if }i\neq j\wedge\mathbf{v}_{i}\in\text{NN}_{k}(\mathbf{v}_{j})\\ 0,&\textrm{otherwise}\end{cases}$ (2)
@@ 5.1 Implementation details
Figure 5: Embeddings retain ordered relations across domains (t-SNE visualization).
During the Deployment stage, we pick the 1% and 99% percentile values of each target variable as the lower and upper bound and split the entire dataset into 5 groups. We implement MDR in a few-shot setup. We use standard hyperparameters that are reported to work well across different setups ($\gamma=3$, $\alpha=0.99$) [16], and use the entire support set for diffusion, setting $k=N$ and assigning the final values with $k_{v}=2N$, with the effects of $k_{v}$ shown in Supplementary Fig. 6. We refer to this approach as GOL+MDR.
@@ 5.3 Quantitative analysis
We present a comparison of our approach with common baselines and state-of-the-art methods in Table 2. We compare against regression on the target domain without adaptation (Regression), regression on the target domain with linear calibration using the support set (Regression cal.), weighted k-nearest neighbor prediction from support examples (kNN), fine-tuning with the support set (FT) [25], geometric order learning with weighted kNN prediction (GOL) [24].
@@
Source Denmark France
Target Spain Slovakia Slovenia Spain Slovakia Slovenia avg
Height Regression 0.35 -.05 0.23 0.61 0.11 0.13 0.23
Regression cal. $0.17\scriptstyle{\pm 0.12}$ $0.49\scriptstyle{\pm 0.10}$ $0.67\scriptstyle{\pm 0.03}$ $0.61\scriptstyle{\pm 0.06}$ $0.42\scriptstyle{\pm 0.06}$ $0.67\scriptstyle{\pm 0.02}$ 0.50
FT [25] $0.34\scriptstyle{\pm 0.00}$ $0.45\scriptstyle{\pm 0.02}$ $0.60\scriptstyle{\pm 0.02}$ $0.69\scriptstyle{\pm 0.01}$ $\mathbf{0.49\scriptstyle{\pm 0.00}}$ $0.50\scriptstyle{\pm 0.09}$ $0.51$
kNN $0.21\scriptstyle{\pm 0.06}$ $0.45\scriptstyle{\pm 0.04}$ $0.68\scriptstyle{\pm 0.03}$ $0.55\scriptstyle{\pm 0.04}$ $0.38\scriptstyle{\pm 0.06}$ $0.63\scriptstyle{\pm 0.03}$ 0.48
GOL [24] $0.21\scriptstyle{\pm 0.12}$ $0.35\scriptstyle{\pm 0.13}$ $0.67\scriptstyle{\pm 0.04}$ $0.62\scriptstyle{\pm 0.05}$ $0.41\scriptstyle{\pm 0.08}$ $0.65\scriptstyle{\pm 0.03}$ $0.49$
GOL+MDR $\mathbf{0.60\scriptstyle{\pm 0.04}}$ $0.49\scriptstyle{\pm 0.06}$ $0.65\scriptstyle{\pm 0.05}$ $0.65\scriptstyle{\pm 0.04}$ $0.46\scriptstyle{\pm 0.05}$ $0.66\scriptstyle{\pm 0.04}$ $0.59$
GOL+FT+MDR $0.56\scriptstyle{\pm 0.07}$ $\mathbf{0.57\scriptstyle{\pm 0.03}}$ $\mathbf{0.69\scriptstyle{\pm 0.01}}$ $\mathbf{0.70\scriptstyle{\pm 0.01}}$ $0.47\scriptstyle{\pm 0.01}$ $\mathbf{0.71\scriptstyle{\pm 0.01}}$ $\mathbf{0.62}$
Count Regression $\mathbf{0.53}$ 0.36 0.43 0.51 0.43 0.37 0.44
Regression cal. $0.51\scriptstyle{\pm 0.04}$ $0.53\scriptstyle{\pm 0.05}$ $0.59\scriptstyle{\pm 0.03}$ $0.53\scriptstyle{\pm 0.07}$ $0.61\scriptstyle{\pm 0.04}$ $\mathbf{0.58\scriptstyle{\pm 0.04}}$ $\mathbf{0.56}$
FT [25] $\mathbf{0.53\scriptstyle{\pm 0.00}}$ $0.48\scriptstyle{\pm 0.00}$ $\mathbf{0.60\scriptstyle{\pm 0.01}}$ $\mathbf{0.54\scriptstyle{\pm 0.00}}$ $0.53\scriptstyle{\pm 0.01}$ $0.40\scriptstyle{\pm 0.00}$ $0.51$
kNN $0.42\scriptstyle{\pm 0.07}$ $0.43\scriptstyle{\pm 0.06}$ $0.55\scriptstyle{\pm 0.03}$ $0.51\scriptstyle{\pm 0.06}$ $0.55\scriptstyle{\pm 0.14}$ $0.55\scriptstyle{\pm 0.08}$ 0.50
GOL [24] $0.51\scriptstyle{\pm 0.03}$ $\mathbf{0.56\scriptstyle{\pm 0.03}}$ $0.58\scriptstyle{\pm 0.05}$ $0.51\scriptstyle{\pm 0.05}$ $0.58\scriptstyle{\pm 0.05}$ $0.46\scriptstyle{\pm 0.06}$ $0.53$
GOL+MDR $0.51\scriptstyle{\pm 0.04}$ $0.55\scriptstyle{\pm 0.06}$ $0.59\scriptstyle{\pm 0.03}$ $\mathbf{0.54\scriptstyle{\pm 0.04}}$ $0.61\scriptstyle{\pm 0.03}$ $0.54\scriptstyle{\pm 0.04}$ $\mathbf{0.56}$
GOL+FT+MDR $\mathbf{0.53\scriptstyle{\pm 0.01}}$ $0.49\scriptstyle{\pm 0.10}$ $0.59\scriptstyle{\pm 0.05}$ $0.51\scriptstyle{\pm 0.09}$ $\mathbf{0.63\scriptstyle{\pm 0.01}}$ $0.53\scriptstyle{\pm 0.03}$ $0.55$
Cover Regression 0.11 0.70 $\mathbf{0.81}$ 0.80 0.77 0.76 0.66
Regression cal. $0.41\scriptstyle{\pm 0.04}$ $0.75\scriptstyle{\pm 0.01}$ $0.77\scriptstyle{\pm 0.04}$ $0.83\scriptstyle{\pm 0.02}$ $0.79\scriptstyle{\pm 0.01}$ $0.77\scriptstyle{\pm 0.04}$ $\mathbf{0.72}$
FT [25] $0.27\scriptstyle{\pm 0.00}$ $0.44\scriptstyle{\pm 0.22}$ $0.75\scriptstyle{\pm 0.01}$ $0.81\scriptstyle{\pm 0.01}$ $0.81\scriptstyle{\pm 0.00}$ $\mathbf{0.81\scriptstyle{\pm 0.00}}$ $0.65$
kNN $0.47\scriptstyle{\pm 0.04}$ $0.69\scriptstyle{\pm 0.02}$ $0.71\scriptstyle{\pm 0.06}$ $0.75\scriptstyle{\pm 0.04}$ $0.73\scriptstyle{\pm 0.05}$ $0.68\scriptstyle{\pm 0.07}$ 0.67
GOL [24] $0.51\scriptstyle{\pm 0.06}$ $\mathbf{0.76\scriptstyle{\pm 0.02}}$ $0.69\scriptstyle{\pm 0.10}$ $0.81\scriptstyle{\pm 0.03}$ $\mathbf{0.82\scriptstyle{\pm 0.02}}$ $0.71\scriptstyle{\pm 0.09}$ $\mathbf{0.72}$
GOL+MDR $0.57\scriptstyle{\pm 0.08}$ $0.71\scriptstyle{\pm 0.03}$ $0.70\scriptstyle{\pm 0.06}$ $0.82\scriptstyle{\pm 0.03}$ $0.79\scriptstyle{\pm 0.03}$ $0.70\scriptstyle{\pm 0.06}$ $0.71$
GOL+FT+MDR $\mathbf{0.65\scriptstyle{\pm 0.05}}$ $0.74\scriptstyle{\pm 0.04}$ $0.64\scriptstyle{\pm 0.05}$ $\mathbf{0.84\scriptstyle{\pm 0.01}}$ $0.76\scriptstyle{\pm 0.03}$ $0.70\scriptstyle{\pm 0.07}$ $\mathbf{0.72}$
Figure 7: Transductive vs. inductive performance across different setups and tasks. Pairwise performance differences between inductive and transductive methods evolve with off-the-shelf regression performance drop (source domain performance - target domain performance), an ad-hoc indicator of domain gap.
@@
We note that the performance of direct cross-domain regression can serve as an indicator of a multifactorial domain gap. For example, on cover regression, where the label shift is limited, direct regression performs well, except for Denmark $\rightarrow$ Spain which is arguably the largest visual and semantic domain shift (Supplementary Fig. 2). We plot in Figure 7 a pairwise comparison of the two best-performing inductive and transductive methods against off-the-shelf regression performance drop. When the implied domain gap becomes large enough for inductive methods to fail, transductive methods are favored.
We repeat experiments with VGG16 [43], another popular feature extractor. GOL+MDR and GOL outperformed regression baselines for height and count with a larger margin than using the ViT-B/16 backbone (Supplementary Table 2).

### v2

@@ Abstract
Image-level regression is an important task in Earth observation, where visual domain and label shifts are a core challenge hampering generalization. However, cross-domain regression within remote sensing data remains understudied due to the absence of suited datasets. We introduce a new dataset with aerial and satellite imagery in five countries with three forest-related regression tasks [^2] Dataset and code available here: dgominski.github.io/drift/. To match real-world applicative interests, we compare methods through a restrictive setup where no prior on the target domain is available during training, and models are adapted with limited information during testing. Building on the assumption that ordered relationships generalize better, we propose manifold diffusion for regression as a strong baseline for transduction in low-data regimes. Our comparison highlights the comparative advantages of inductive and transductive methods in cross-domain regression.
@@ 1 Introduction
Data availability and downstream applications define the scope of models (regional [36] – national [45] – continental [5]), but they don’t necessarily match the visual and semantic domains defined by biomes and environmental conditions. Intuitively, a model trained on a dataset should perform relatively well on datasets that are spatially close and visually similar, although it remains unclear to what extent.
To better assess generalization performance and go towards universal models that can be trained in countries with high data availability and applied to any country without prior knowledge or retraining, we introduce the DRIFT dataset for universal, low-data domain-adaptive regression.
Building on recent advances in order learning, we hypothesize that order relationships generalize well across domains (Fig. 1). We enforce a well-ordered embedding space in the source domain, then adapt to target domains in low-data setups. Notably, this strong regularization during training does not hamper source-domain performance and creates a clean embedding space to propagate similarities at test-time, which opens the way for advanced low-data inference in the form of manifold diffusion. Our contributions are as follows:
@@
  • We compare inductive and transductive methods across datasets and tasks, in a low-shot source-free domain adaptation setup. We show that transductive methods have an advantage when the domain gap becomes predominant.
@@ 2.3 Image-Level Regression in Earth Observation
The increasing availability of vast and diverse collections of remote sensing imagery and datasets presents both opportunities and challenges for computer vision tasks. To date, most works utilizing remote sensing imagery focus on image classification and semantic segmentation [2, 53, 4]. For example, to count trees, some works rely on the detection of individual trees [33, 5], but manual annotation relies on expert knowledge and is extremely time-consuming. Given the variations in ground resolution, image quality, and geographic areas, small objects and occlusion phenomena can make exhaustive annotation virtually impossible. Other labels on woody biomass or canopy height cannot be manually labeled. Therefore, labels often have to be sourced from field measures [37, 9], or derived from low-resolution dedicated sensors such as GEDI [28, 14], and remain noisy due to the varying viewing angles, technical inaccuracies in positioning systems [17] or temporal mismatch [59, 6]. As such, accurate labels are often found at a lower resolution than the ground resolution of the image. This makes image-level regression a core task in computer vision for Earth observation, as it offers a way to directly predict the information at the available resolution [12, 7]. To our knowledge, there is no existing dataset for cross-domain image-level regression using sub-meter resolution remote sensing imagery (Supplementary Tables 2 and 3).
@@ 3 DRIFT: Domain-Adaptive Regression for Forest Monitoring across Countries
    Tree count: number of overstory (visible from an overhead perspective) trees in the images.
@@
Dataset Collection. We built the dataset from national aerial orthophotography with various ground sampling distances (GSD) for Denmark (DK), Spain (SP), France (FR), and Slovakia (SK). All aerial images contain three spectral channels: Red (R), Green (G), and Blue (B). For Slovenia (SI), we used very high-resolution satellite imagery from commercial nanosatellites (SkySat). We acquired the orthorectified surface reflectance products at a spatial resolution of 50cm. Characteristics of each subset in the DRFIT are summarized in Table 1. Refer to Supplementary section 1 for extended details regarding the dataset curation, data origins, and visual examples.
Table 1: DRIFT Dataset characteristics. GSD and Height in meters, Cover in percents.
@@
Figure 2: Challenging examples in the DRIFT dataset: despite similar visual content, images can have different values in the label space. More visual examples can be found in Supplementary Fig. 1.
@@
Figure 3: Label shift varies across countries and tasks. We plot Wasserstein distances between label distributions in country subsets to indicate the level of label shift.
Regression Tasks. Canopy height, tree count, and tree cover heavily depend on tree species, local climate, and landscape characteristics. Height prediction from optical images is inherently challenging due to limited visual clues about depth from an aerial viewpoint, especially if the ground is not visible, but it has been shown possible using satellite imagery [28, 41]. Counts depend on the model’s interpretation of individual crowns, which is subject to extreme variations even within a species [22]. Cover fraction is intuitively the easiest, as the model only needs to separate trees from background pixels. The difficulty depends on the level of similarity between short vegetation ($<5$m) and trees.
@@ 4 Method
We explore the potential of ordered embeddings and simple transductive learning on the cross-domain challenges of the DRIFT dataset (see Fig. 4). We train the embedding extractor on the source domain, using geometric order learning to enforce structure. We extract embeddings on the target domain, and perform adaptation for regression using a few labels. Optionally, we refine predictions to take into account local and global manifold structure with diffusion.
@@ 4.1 Preliminaries
Let $\mathcal{D}_{s}=\{(\textbf{x}_{i}^{s},y_{i}^{s})\}$ and $\mathcal{D}_{t}=\{(\textbf{x}_{i}^{t},y_{i}^{t})\}$ denote the source and target domains respectively. $y^{s}$ and $y^{t}$ take continuous values in , and we note that their distributions are different. Notably, it is common in vegetation mapping that their support $[y^{s}_{min},y^{s}_{max}]$ and $[y^{t}_{min},y^{t}_{max}]$ differs. For example, $y^{s}_{max}\neq y^{t}_{max}$ in height mapping due to different species.
@@
A simple baseline method would be to train a regressor $f(\textbf{x})=y$ on $\mathcal{D}_{s}$ and apply it to $\mathcal{D}_{t}$. Due to domain shifts, this performs generally poorly. Instead, we consider a feature extractor $h(\textbf{x})=\mathbf{v}$, with $\mathbf{v}\in^{d}$ and $\lVert\mathbf{v}\lVert_{2}=1$, and frame the problem as cross-domain transduction: $y^{t}$ is inferred from the relationships between the embeddings in $\mathcal{D}_{t}$.
@@ 4.2 Manifold Diffusion for Regression
We start by introducing the diffusion framework originally proposed in [67]. Assuming a set $X=\{\textbf{x}_{1},\textbf{x}_{2},...,\textbf{x}_{n}\}$ and $\mathbf{v}_{i}=h(\textbf{x}_{i})$, the affinity matrix $W\in^{n\times n}$ is constructed by calculating the pairwise cosine similarity between samples:
$W_{ij}:=\begin{cases}[\mathbf{v}_{i}^{\top}\mathbf{v}_{j}]_{+}^{\gamma},&\textrm{if }i\neq j\wedge\mathbf{v}_{i}\in\text{NN}_{k}(\mathbf{v}_{j})\\ 0,&\textrm{otherwise}\end{cases}$ (2)
@@ 5.1 Implementation details
Figure 5: Embeddings retain ordered relations across domains (t-SNE visualization). From left to right: source domain training data, source domain testing data, and three target domains, respectively.
During the Deployment stage, we pick the 1% and 99% percentile values of each target variable as the lower and upper bound and split the entire dataset into 5 groups.
We implement MDR in a few-shot setup. We use standard hyperparameters that are reported to work well across different setups ($\gamma=3$, $\alpha=0.99$) [20], and use the entire support set for diffusion, setting $k=N$ and assigning the final values with $k_{v}=2N$, with the effects of $k_{v}$ shown in Supplementary Fig. 6. We refer to this approach as GOL+MDR.
@@ 5.3 Quantitative Analysis
We compare our approach with common baselines and state-of-the-art methods in Table 2. We compare against regression on the target domain without adaptation (Regression), regression on the target domain with linear calibration using the support set (Regression cal.), weighted k-nearest neighbor prediction from support examples (kNN), fine-tuning with the support set (FT) [31], geometric order learning with weighted kNN prediction (GOL) [30].
@@
Source Denmark France Target Spain Slovakia Slovenia Spain Slovakia Slovenia avg Height Regression 0.35 -.05 0.23 0.61 0.11 0.13 0.23 Regression cal. $0.17\scriptstyle{\pm 0.12}$ $0.49\scriptstyle{\pm 0.10}$ $0.67\scriptstyle{\pm 0.03}$ $0.61\scriptstyle{\pm 0.06}$ $0.42\scriptstyle{\pm 0.06}$ $0.67\scriptstyle{\pm 0.02}$ 0.50 FT [31] $0.34\scriptstyle{\pm 0.00}$ $0.45\scriptstyle{\pm 0.02}$ $0.60\scriptstyle{\pm 0.02}$ $0.69\scriptstyle{\pm 0.01}$ $\mathbf{0.49\scriptstyle{\pm 0.00}}$ $0.50\scriptstyle{\pm 0.09}$ $0.51$ kNN $0.21\scriptstyle{\pm 0.06}$ $0.45\scriptstyle{\pm 0.04}$ $0.68\scriptstyle{\pm 0.03}$ $0.55\scriptstyle{\pm 0.04}$ $0.38\scriptstyle{\pm 0.06}$ $0.63\scriptstyle{\pm 0.03}$ 0.48 GOL [30] $0.21\scriptstyle{\pm 0.12}$ $0.35\scriptstyle{\pm 0.13}$ $0.67\scriptstyle{\pm 0.04}$ $0.62\scriptstyle{\pm 0.05}$ $0.41\scriptstyle{\pm 0.08}$ $0.65\scriptstyle{\pm 0.03}$ $0.49$ GOL+MDR $\mathbf{0.60\scriptstyle{\pm 0.04}}$ $0.49\scriptstyle{\pm 0.06}$ $0.65\scriptstyle{\pm 0.05}$ $0.65\scriptstyle{\pm 0.04}$ $0.46\scriptstyle{\pm 0.05}$ $0.66\scriptstyle{\pm 0.04}$ $0.59$ GOL+FT+MDR $0.56\scriptstyle{\pm 0.07}$ $\mathbf{0.57\scriptstyle{\pm 0.03}}$ $\mathbf{0.69\scriptstyle{\pm 0.01}}$ $\mathbf{0.70\scriptstyle{\pm 0.01}}$ $0.47\scriptstyle{\pm 0.01}$ $\mathbf{0.71\scriptstyle{\pm 0.01}}$ $\mathbf{0.62}$ Count Regression $\mathbf{0.53}$ 0.36 0.43 0.51 0.43 0.37 0.44 Regression cal. $0.51\scriptstyle{\pm 0.04}$ $0.53\scriptstyle{\pm 0.05}$ $0.59\scriptstyle{\pm 0.03}$ $0.53\scriptstyle{\pm 0.07}$ $0.61\scriptstyle{\pm 0.04}$ $\mathbf{0.58\scriptstyle{\pm 0.04}}$ $\mathbf{0.56}$ FT [31] $\mathbf{0.53\scriptstyle{\pm 0.00}}$ $0.48\scriptstyle{\pm 0.00}$ $\mathbf{0.60\scriptstyle{\pm 0.01}}$ $\mathbf{0.54\scriptstyle{\pm 0.00}}$ $0.53\scriptstyle{\pm 0.01}$ $0.40\scriptstyle{\pm 0.00}$ $0.51$ kNN $0.42\scriptstyle{\pm 0.07}$ $0.43\scriptstyle{\pm 0.06}$ $0.55\scriptstyle{\pm 0.03}$ $0.51\scriptstyle{\pm 0.06}$ $0.55\scriptstyle{\pm 0.14}$ $0.55\scriptstyle{\pm 0.08}$ 0.50 GOL [30] $0.51\scriptstyle{\pm 0.03}$ $\mathbf{0.56\scriptstyle{\pm 0.03}}$ $0.58\scriptstyle{\pm 0.05}$ $0.51\scriptstyle{\pm 0.05}$ $0.58\scriptstyle{\pm 0.05}$ $0.46\scriptstyle{\pm 0.06}$ $0.53$ GOL+MDR $0.51\scriptstyle{\pm 0.04}$ $0.55\scriptstyle{\pm 0.06}$ $0.59\scriptstyle{\pm 0.03}$ $\mathbf{0.54\scriptstyle{\pm 0.04}}$ $0.61\scriptstyle{\pm 0.03}$ $0.54\scriptstyle{\pm 0.04}$ $\mathbf{0.56}$ GOL+FT+MDR $\mathbf{0.53\scriptstyle{\pm 0.01}}$ $0.49\scriptstyle{\pm 0.10}$ $0.59\scriptstyle{\pm 0.05}$ $0.51\scriptstyle{\pm 0.09}$ $\mathbf{0.63\scriptstyle{\pm 0.01}}$ $0.53\scriptstyle{\pm 0.03}$ $0.55$ Cover Regression 0.11 0.70 $\mathbf{0.81}$ 0.80 0.77 0.76 0.66 Regression cal. $0.41\scriptstyle{\pm 0.04}$ $0.75\scriptstyle{\pm 0.01}$ $0.77\scriptstyle{\pm 0.04}$ $0.83\scriptstyle{\pm 0.02}$ $0.79\scriptstyle{\pm 0.01}$ $0.77\scriptstyle{\pm 0.04}$ $\mathbf{0.72}$ FT [31] $0.27\scriptstyle{\pm 0.00}$ $0.44\scriptstyle{\pm 0.22}$ $0.75\scriptstyle{\pm 0.01}$ $0.81\scriptstyle{\pm 0.01}$ $0.81\scriptstyle{\pm 0.00}$ $\mathbf{0.81\scriptstyle{\pm 0.00}}$ $0.65$ kNN $0.47\scriptstyle{\pm 0.04}$ $0.69\scriptstyle{\pm 0.02}$ $0.71\scriptstyle{\pm 0.06}$ $0.75\scriptstyle{\pm 0.04}$ $0.73\scriptstyle{\pm 0.05}$ $0.68\scriptstyle{\pm 0.07}$ 0.67 GOL [30] $0.51\scriptstyle{\pm 0.06}$ $\mathbf{0.76\scriptstyle{\pm 0.02}}$ $0.69\scriptstyle{\pm 0.10}$ $0.81\scriptstyle{\pm 0.03}$ $\mathbf{0.82\scriptstyle{\pm 0.02}}$ $0.71\scriptstyle{\pm 0.09}$ $\mathbf{0.72}$ GOL+MDR $0.57\scriptstyle{\pm 0.08}$ $0.71\scriptstyle{\pm 0.03}$ $0.70\scriptstyle{\pm 0.06}$ $0.82\scriptstyle{\pm 0.03}$ $0.79\scriptstyle{\pm 0.03}$ $0.70\scriptstyle{\pm 0.06}$ $0.71$ GOL+FT+MDR $\mathbf{0.65\scriptstyle{\pm 0.05}}$ $0.74\scriptstyle{\pm 0.04}$ $0.64\scriptstyle{\pm 0.05}$ $\mathbf{0.84\scriptstyle{\pm 0.01}}$ $0.76\scriptstyle{\pm 0.03}$ $0.70\scriptstyle{\pm 0.07}$ $\mathbf{0.72}$
@@
Figure 7: Transductive methods outperform inductive methods when the domain gap is larger. We plot pairwise performance differences between inductive and transductive methods on the y-axis, with off-the-shelf regression performance drop (source domain performance minus target domain performance) on the x-axis, an ad-hoc indicator of domain gap.
@@
We note that the performance of direct cross-domain regression can serve as an indicator of a multifactorial domain gap. For example, on cover regression, where the label shift is limited, direct regression performs well, except for Denmark $\rightarrow$ Spain which is arguably the largest visual and semantic domain shift (Supplementary Fig. 2). We plot in Fig. 7 a pairwise comparison of the two best-performing inductive and transductive methods against off-the-shelf regression performance drop. When the implied domain gap becomes large enough for inductive methods to fail, transductive methods are favored.
We repeat experiments with VGG16 [57], another popular feature extractor. GOL+MDR and GOL outperformed regression baselines for height and count with larger margins than using the ViT-B/16 backbone (Supplementary Table 4).